a number is such that thrice it is 14 less than 5times the number find the no
Answer by macston(5194)

number=N
3N=5N-14 subtract 3N from each side
3N-3N=5N-14-3N
0=2N-14 add 14 to each side
0+14=2N-14+14
14=2N divide each side by 2
7=N
CHECK
3(7)=5(7)-14
21=35-14
21=21